# Between Circuits and Chomsky: Pre-pretraining on Formal Languages Imparts Linguistic Biases

### v1

@@ 8 Discussion
Our results suggest that pre-pretraining on formal languages with hierarchical dependencies representable in C-RASP can improve the downstream loss and linguistic generalization of transformers. As we conjectured in the introduction, context-sensitivity alone is not sufficient for positive transfer. The copy language $ww$ contains cross-serial dependencies, but it contains no notion of hierarchy and is not definable in C-RASP, and pre-pretraining on it hurts natural language performance. Conversely, $k$-Dyck and $k$-Shuffle Dyck both model hierarchical dependencies and confer a better inductive bias than natural language. $k$-Shuffle Dyck, the language implementable in C-RASP of the two, leads to the largest pre-pretraining gains.
As shown in §5, learning structure also has downstream effects. The main subnetwork identified in pre-pretraining does not contain the attention heads important for morphological acceptability judgments. Nevertheless, pre-pretraining does have a positive effect on models learning morphology. Thus, pre-pretraining may have the potential to improve models beyond their performance on hierarchical, rule-based tasks. At the same time, although the overall trend is positive, performance does decrease on some of the individual BLiMP tasks. So pre-pretraining has downstream tradeoffs as well.
We found pre-pretraining on formal languages to be beneficial even for the 1B parameter training setting. Furthermore, the MRS between formal and natural language is greater than one (Table 5), meaning that one token of formal language in pre-pretraining substitutes for more than one token of natural language in pretraining. This is a surprising result from the perspective of statistical learning theory Vapnik (2000), in that we observe faster convergence by swapping in data from a different task. Initializations affect learning dynamics McCoy et al. (2020); Sellam et al. (2022), and the better initialization from pre-pretraining materially speeds up natural language pretraining.
@@ 9 Limitations
In this work, we considered blocked training, where we first train on formal language followed by natural language. We chose blocked training because a good initialization from formal language can be easily plugged into existing pretraining pipelines, but mixing formal and natural language during training could lead to better performance Korbak et al. (2023). We also evaluated efficiency in a setting where pretraining data is plentiful, and one can train without running several epochs over the same data. A low-resource setting would also be interesting, and may yield different scaling properties with respect to pre-pretraining data Muennighoff et al. (2023). This would be particularly relevant for non-English languages Samuel et al. (2023); Martin et al. (2020). Finally, our work only considers transformers. Circuit complexity has also quantified the expressive power of neural networks like RNNs and state-space models Merrill et al. (2020, 2024), and it would be interesting to extend our results to these architectures.
@@ Appendix B Hyperparameters
Table 3: BLiMP Warstadt et al. (2020a) dataset examples showing grammatical (positive) and ungrammatical (negative) sentence pairs
@@
Table 4: Verbatim in-context retrieval Armeni et al. (2022, 2024) examples. A good language model should recognize from context that the list is repeated.
Figure 9: Grammaticality judgment (BLiMP) accuracies by task.
@@
Table 5: Evaluating models at the optimal amount of pre-pretraining $t_{0}^{*}$ for each formal language. “Documents $\downarrow$” is the proportion of documents in the C4 validation set with lower loss than the model trained without pre-pretraining. 1-Dyck, $k$-Dyck, and $k$-Shuffle-Dyck all have marginal rates of substitution (MRS) greater than 1, indicating that pre-pretraining is more efficient than not pre-pretraining. $k$-Shuffle-Dyck performs the best overall on our evaluation metrics.
Code 1: Implementation of a $k$-Shuffle Dyck sequence generator.

### v2

@@ 8 Discussion
We have found that pre-pretraining on formal languages can improve the language modeling loss and linguistic generalization abilities of transformer language models. In fact, pre-pretraining on some formal languages was more effective than increasing the amount of natural language training data by the same amount: the inductive bias conferred by the formal language was more helpful than additional in-distribution data. While most of the experiments in this paper were with 160M-parameter models, we also found benefits from pre-pretraining in the 1B-parameter setting.
We hypothesized that the languages that are most effective in this paradigm are those that, first, feature hierarchical dependencies, and second, are representable in C-RASP, and therefore readily learnable by transformers. The first part of this hypothesis is supposed by the superior performance of $k$-Dyck and $k$-Shuffle Dyck, the languages with hierarchical dependencies, relative to other languages. Our evidence for the importance of expressibility in C-RASP is less clear: $k$-Shuffle Dyck clearly outperformed $k$-Dyck, which is not expressible in C-RASP, on one of the evaluations, and also required fewer steps of pre-pretraining than the other languages (§4). While natural language is arguably context-sensitive, in the Chomsky hierarchy sense, not every context-sensitive language was beneficial in pre-pretraining: in fact, pre-pretraining was harmful when we used the copy language $ww$, which, while context-sensitive since it contains cross-serial dependencies, does not illustrate the notion of hierarchy and is not definable in C-RASP. That being said, since $k$-Dyck performed almost as well as $k$-Shuffle Dyck, there may exist a sharper characterization of the class of formal languages that confer a helpful inductive bias than defined by our expressivity hypothesis; experiments with a larger sample of formal languages would be needed to progress towards such a characterization.
The marginal rate of substitution between formal and natural language is greater than one (Table 5), meaning that one token of formal language in pre-pretraining substitutes for more than one token of natural language in pretraining. This is a surprising result from the perspective of statistical learning theory Vapnik (2000), in that we observe faster convergence by swapping in data from a different distribution. We hypothesize that initialization can have a critical effect on learning dynamics McCoy et al. (2020b); Sellam et al. (2022), and pre-pretraining on formal language produces a initialization that is favorable to natural language learning.
@@ 9 Limitations
In this work, we considered blocked training, where we first train on formal language and then on natural language. While blocked training has the advantage that the initialization produced by formal language pre-pretraining can then be distributed and easily plugged into existing pretraining pipelines, it is possible that the optimal training regimen involves mixing formal and natural language during training Korbak et al. (2023). We also evaluated the effectiveness of pre-pretraining in a setting where natural language pretraining data is plentiful, as it is for English, such that it is possible to train the model for a considerable number of tokens without processing the same data multiple times over several epochs. We hypothesize that pre-pretraining will be even more effective for low-resource natural languages, and may yield different scaling properties with respect to pre-pretraining data Muennighoff et al. (2023). Relatedly, a natural extension to this project is establishing scaling laws for pre-pretraining; the benefit of pre-pretraining beyond 1 billion parameters and 1.6 billion tokens is currently unknown. Finally, our work only considers transformers. Circuit complexity has also quantified the expressive power of neural networks like RNNs and state-space models Merrill et al. (2020, 2024), and it would be interesting to extend our results to these architectures.
@@ Appendix B Hyperparameters
Table 3: Examples from the BLiMP dataset Warstadt et al. (2020a): matched pairs of grammatical (positive, left) and ungrammatical (negative, right) sentences. We expect the language model to assign a higher probability to the grammatical sentence in each pair.
@@
Table 4: Verbatim in-context retrieval Armeni et al. (2022, 2024) examples. We expect a good language model to recognize based on the context that the list is repeated, retrieve the appropriate items from the first repetition of the list, and assign these items a very high probability.
Figure 8: Accuracy on BLiMP by grammatical phenomenon. The full names of the phenomena are: anaphor agreement, argument structure, binding, control/raising, determiner-noun agreement, ellipsis, filler-gap dependencies, irregular forms, island effects, negative polarity item licensing, quantifiers, and subject-verb agreement.
Figure 9: Performance changes on BLiMP after pruning half the attention heads from the model trained on $k$-Shuffle Dyck (see §5). The largest declines are on subject-verb agreement, irregular forms, and anaphor agreement. These categories require knowledge about word forms, and the sentences within these categories are generally simple (around 4 words).
@@
Table 5: Evaluating models at the optimal amount of pre-pretraining $t_{0}^{*}$ for each formal language (see §4). “Documents $\downarrow$” is the proportion of documents in the C4 validation set where the model has a lower loss than the model trained without pre-pretraining. 1-Dyck, $k$-Dyck, and $k$-Shuffle-Dyck all have marginal rates of substitution (MRS) greater than 1, indicating that pre-pretraining is more efficient than not pre-pretraining. $k$-Shuffle-Dyck performs the best overall on our evaluation metrics.
Figure 10: Implementation of a $k$-Shuffle Dyck sequence generator.